Your cap size is based on the circumference of your head (in inches). For a certain range of head circumferences c, cap size s can be modeled by the following equation:
s=c-1/3
If you wear a size 7 cap, what is your head circumference? Please show all of your work.
thank you!

Answer by stanbon(75887)

cap size s can be modeled by the following equation:
s=(c-1)/3
If you wear a size 7 cap, what is your head circumference?
-----------------------
size = (circumference - 1)/3
-------
7 = (c-1)/3
c-1 = 21
c = 22 cm
=========================================
Cheers,
Stan H.